OpenStudy (anonymous):

I WILL MEDAL AND FAN!! HELP ME UNDERSTAND Choose the correct description of the graph of the compound inequality: 3x + 1 > 7 and 4x less than or equal to 24 A number line with an open circle on 2, a closed circle on 6, and shading in between A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right A number line with a closed circle on 2, an open circle on 6, and shading in between

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi again

OpenStudy (misty1212):

can you solve \[3x+1>7\]? takes two steps only

OpenStudy (anonymous):

subtract 1, then you get 3x>6, but you divide by 3 to get x>2

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

so its A or b, not c and d

OpenStudy (misty1212):

so one inequality is \(x>2\) lets solve the other one \[4x\leq 24\]

OpenStudy (anonymous):

i think that its b?

OpenStudy (misty1212):

idk lets solve the inequality , then we will know without guessing there is more work to do

OpenStudy (anonymous):

wait i see what i di wrong, i mixed up the answers. i think its a, but we shuld solve anyway.

OpenStudy (misty1212):

takes one step to solve \[4x\leq 28\]

OpenStudy (misty1212):

oops i meant \[4x\leq 24\] what do you get ?

OpenStudy (anonymous):

x<=6

OpenStudy (anonymous):

so A

OpenStudy (misty1212):

ok so it is \[2<x\leq 6\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]